OpenStudy (anonymous):

solve the integral: \[\int\limits_{0}^{\infty}((arctanx)/(1+x))dx\]

OpenStudy (mr.math):

Is it 1+x or 1+x^2?

OpenStudy (anonymous):

1 + x^2 sorry for that mistake.

OpenStudy (anonymous):

I don't think it converges

OpenStudy (mr.math):

This is very easy. Just substitute \(\tan^{-1}(x)\). The integral will be \(\frac{1}{2}(\tan^{-1}x)^+c\).

OpenStudy (mr.math):

\[\frac{1}{2}\arctan^{2}x+c\]

OpenStudy (anonymous):

ok gimme a sec, I'm doing it on paper

OpenStudy (anonymous):

OK got it, thanks

OpenStudy (mr.math):

But your integral is definite. So it would be \[\frac{1}{2}\arctan^{2}(x)|_0^{\infty}.\]

OpenStudy (mr.math):

You're welcome.

OpenStudy (anonymous):

Oh Cool, I thought 1 + x